# Group-in-Group Policy Optimization for LLM Agent Training

### v1

@@ Abstract
Recent advances in group-based reinforcement learning (RL) have driven frontier large language models (LLMs) in single-turn tasks like mathematical reasoning. However, their scalability to long-horizon LLM agent training remains limited. Unlike static tasks, agent-environment interactions unfold over many steps and often yield sparse or delayed rewards, making credit assignment across individual steps significantly more challenging. In this work, we propose Group-in-Group Policy Optimization (GiGPO), a novel RL algorithm that achieves fine-grained credit assignment for LLM agents while preserving the appealing properties of group-based RL: critic-free, low memory, and stable convergence. GiGPO introduces a two-level structure for estimating relative advantage: (i) At the episode-level, GiGPO computes macro relative advantages based on groups of complete trajectories; (ii) At the step-level, GiGPO introduces an anchor state grouping mechanism that retroactively constructs step-level groups by identifying repeated environment states across trajectories. Actions stemming from the same state are grouped together, enabling micro relative advantage estimation. This hierarchical structure effectively captures both global trajectory quality and local step effectiveness without relying on auxiliary models or additional rollouts. We evaluate GiGPO on two challenging agent benchmarks, ALFWorld and WebShop, using Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct. Crucially, GiGPO delivers fine-grained per-step credit signals and achieves performance gains of > 12% on ALFWorld and > 9% on WebShop over the GRPO baseline: all while maintaining the same GPU memory overhead, identical LLM rollout, and incurring little to no additional time cost.††footnotetext: Code: https://github.com/langfengQ/verl-agent
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3, 4] have leapt from static question-answer systems to versatile agents that perceive, reason, and act in open-ended environments. For instance, they now power embodied assistants that navigate simulated homes [5, 6], mobile and web navigators that plan multi-step browsing sessions [7, 8, 9], and autonomous explorers in interactive games [10, 11]. In these settings, LLM agents need to perceive, reason, and act in multi-turn loops, which requires not only language understanding but also long-horizon planning and decision-making.
@@
To address this, we introduce Group-in-Group Policy Optimization (GiGPO), a new group-based RL algorithm that nests two-dimensional notions of credit assignment, better-suited for the multi-step optimization of LLM agents. (i) At the episode level, GiGPO samples a group of complete trajectories under identical task and initial-state conditions, and computes macro relative advantages based on total returns like vanilla GRPO [17]. This captures the overall effectiveness of each trajectory and reflects the completeness of task execution. (ii) At the step level, GiGPO introduces a novel anchor state grouping mechanism for fine-grained relative advantages estimation. Specifically, it retroactively identifies the repeated environment states, termed as anchor states, across trajectories and uses them as anchors to construct step-level groups that allow for localized credit assignment.
@@
We evaluate GiGPO on two long-horizon agent benchmarks: ALFWorld [5], which tests embodied task planning in simulated household environments, and WebShop [21], which simulates complex, goal-driven web interactions. Our experiments with Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [3] show that GiGPO consistently outperforms prompt-based agents, actor-critic baselines, and prior group-based RL methods. In particular, GiGPO injects fine-grained, step-level credit signals that sharpen policy learning of agents over long horizons and achieves performance improvement of > 12% on ALFWorld and > 9% on WebShop over GRPO. These gains come without compromising the core strengths of group-based RL (only < 0.002% time cost), making GiGPO a versatile and high-utility algorithm for the long-horizon optimization of LLM agents.
@@ 2 Related Work
Reinforcement learning for LLM agents. RL has played a pivotal role in enabling LLM agents to operate in dynamic, open-ended environments. Early work applied classical RL algorithms such as DQN [34] to train LLM agents in text-based games [35] and later research [36, 37, 38, 39, 40] started to employ value-based methods, such PPO [41] and AWR [42], in more diverse and interactive agent scenarios including Android device control [43], embodied ALFWorld [5], and card games [44]. More recent approaches have extended RL training to complex web-based and application-centered tasks. For instance, ArCHer [45] and AgentQ [46] target the WebShop benchmark [21], but require intricate designs and computation overhead such as additional value networks or Monte Carlo Tree Search (MCTS) [47]. Going further, LOOP [48] introduces a hybrid method combining REINFORCE leave-one-out (RLOO) [15, 16] with PPO-style updates, achieving state-of-the-art results in AppWorld [49]. RAGEN [50] introduces a trajectory-level GRPO that concatenates all states, intermediate reasoning, and actions into a unified episode-level response. However, it faces scalability challenges in long-horizon tasks (e.g., in ALFWorld, which involves up to 50 steps).
@@ 5 Experiment
In this section, we present empirical evaluations of GiGPO across a variety of LLM agent tasks. Specifically, we aim to demonstrate: (1) the strong ability of GiGPO in training LLM agents; (2) the ablation study of GiGPO; (3) the dynamic trend of step-level group over the course of training; (4) the computational budget of GiGPO.
@@ 5.1 Experiment Setup
Environments. We train the LLM agents on two challenging benchmarks: ALFWorld [5] and WebShop [21]. ALFWorld is an embodied environment designed to assess the ability of LLM agents to perform multi-step decision-making. In each episode, the agent receives a text goal and must accomplish it through multi-turn interaction with the environment. It includes 4,639 task instances across six categories of common household activities: Pick & Place (Pick), Examine in Light (Look), Clean & Place (Clean), Heat & Place (Heat), Cool & Place (Cool), and Pick Two & Place (Pick2). WebShop is a complex, web-based interactive environment designed to test the LLM agents in realistic online shopping scenarios. To complete the task, the agent must interact with a simulated HTML-based shopping website to search for, navigate to, and ultimately purchase a suitable item. It contains over 1.1 million products and 12k user instructions, providing a rich and diverse action space.
Baselines. We compare our approach with a range of competitive baselines: (1) Closed-source LLMs: GPT-4o [1] and Gemini-2.5-Pro [2], which represent state-of-the-art capabilities in general-purpose reasoning and language understanding. (2) Prompting agents: ReAct [28] and Reflexion [29], which rely on in-context prompting to guide multi-step behavior without parameter updates. (3) RL training methods: PPO [41], a widely used actor-critic algorithm that requires an additional value model, and group-based critic-free methods RLOO [15, 16] and GRPO [17], which perform advantage estimation over trajectory groups.
Training details. We use Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [3] as our base models. The weighting coefficient is set to with no further tuning. All RL training methods (including ours and the baselines) use exactly the same hyperparameter configurations. The rollout group size for group-based RL methods is set to 8. The LLM agent is prompted to first generate a chain-of-thought [61] enclosed within <think> </think> tags, followed by the action enclosed within <action> </action> tags. Full training settings and hyperparameter details are provided in Appendix D.1.
5.2 Performance
@@
| Base: Closed-Source Model | ||||||||||
@@ 6 Conclusions and Limitations
In this work, we proposed GiGPO, a novel group-based RL algorithm to tackle the credit assignment challenge in long-horizon LLM agent training. GiGPO introduces a hierarchical advantage estimation that enables fine-grained per-step credit assignment while retaining the efficiency and stability of group-based RL. By retroactively grouping steps that share the same state across trajectories, it achieves this without incurring additional LLM rollout or GPU memory overhead. Empirical evaluations across two complex environments, ALFWorld and WebShop, demonstrate that GiGPO significantly outperforms both prompt-based agents and prior RL methods. A potential limitation of GiGPO is its reliance on state matching for anchor group construction. In highly complex environments, identical states may be hard to detect due to noise or subtle differences. Despite this, GiGPO still retains a strong performance lower bound: in the extreme case where no states are repeated across trajectories (i.e., ), it naturally degrades to GRPO, preserving GRPO’s effectiveness and stability in credit assignment. However, we suggest a more robust solution: introducing state similarity through embeddings or approximate matching, which could better capture structurally equivalent states. We leave this exploration as a promising direction for future work.
@@ Appendix A Source Code
Our code is built upon the veRL framework [64] and extended with several features to support scalable training of long-horizon LLM agents. Key capabilities of our code include: (1) a step-wise multi-turn interaction paradigm that avoids concatenating the entire interaction history into a single input, as done in RAGEN [50], thereby ensuring scalability for very long-horizon optimization, (2) support for parallelized, gym-style group-based environments to enable high-throughput and state-aligned rollouts, (3) a rich suite of environments, including ALFWorld, WebShop, Sokoban, and Gym Cards; and (4) compatibility with both text-based and vision-language agents.
@@ D.1 Details of Training
Computing Details. Experiments using Qwen2.5-1.5B-Instruct are conducted on two NVIDIA H100 GPUs, while those using Qwen2.5-7B-Instruct are trained on four NVIDIA H100 GPUs. Each experiment is trained for a total of 150 training iterations.
@@ D.2 Prompts
The prompts we use for LLM agents are presented in Figure 7 and Figure 8. These prompt templates are constructed using Python-style string formatting, where placeholders enclosed in curly braces ({}) represent semantic slots. These placeholders, such as {task_description}, {step_count}, and {current_observation}, are dynamically populated at runtime via Python’s .format() function. To enrich the agent’s context, we use historical information and set the history length to 2. A fully instantiated example is shown in Figure 9 and Figure 10. This augmented representation is used as the final input to the agent.
The <think>…</think> block instructs the agent to explicitly perform step-by-step reasoning, thereby promoting chain-of-thought style deliberation. The <action>…</action> block is used to clearly indicate the final action decision.

### v2

@@ Abstract
Recent advances in group-based reinforcement learning (RL) have driven frontier large language models (LLMs) in single-turn tasks like mathematical reasoning. However, their scalability to multi-turn LLM agent training remains limited. Unlike static tasks, agent-environment interactions unfold over many steps and often yield sparse or delayed rewards, making credit assignment across individual steps significantly more challenging. In this work, we propose Group-in-Group Policy Optimization (GiGPO), a novel RL algorithm that achieves fine-grained credit assignment for LLM agents while preserving the appealing properties of group-based RL: critic-free, low memory, and stable convergence. GiGPO introduces a two-level structure for estimating relative advantage: (i) At the episode-level, GiGPO computes macro relative advantages based on groups of complete trajectories; (ii) At the step-level, GiGPO introduces an anchor state grouping mechanism that retroactively constructs step-level groups by identifying repeated environment states across trajectories. Actions stemming from the same state are grouped together, enabling micro relative advantage estimation. This hierarchical structure effectively captures both global trajectory quality and local step effectiveness without relying on auxiliary models or additional rollouts. We evaluate GiGPO on challenging agent benchmarks, including ALFWorld and WebShop, as well as tool-integrated reasoning on search-augmented QA tasks, using Qwen2.5-1.5B/3B/7B-Instruct. Crucially, GiGPO delivers fine-grained per-step credit signals, achieves performance gains of > 12% on ALFWorld and > 9% on WebShop over GRPO, and obtains superior performance on QA tasks (42.1% on 3B and 47.2% on 7B): all while maintaining the same GPU memory overhead, identical LLM rollout, and incurring little to no additional time cost. ††footnotetext: Code: https://github.com/langfengQ/verl-agent
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3, 4] have leapt from static question-answer systems to versatile agents that perceive, reason, and act in open-ended environments. For instance, they now power embodied assistants that navigate simulated homes [5, 6], mobile and web navigators that plan multi-step browsing sessions [7, 8, 9, 10], and autonomous explorers in interactive games [11, 12]. In these settings, LLM agents need to perceive, reason, and act in multi-turn loops, which requires not only language understanding but also long-horizon planning and decision-making.
@@
To address this, we introduce Group-in-Group Policy Optimization (GiGPO), a new group-based RL algorithm that nests two-dimensional notions of credit assignment, better-suited for the multi-turn optimization of LLM agents. (i) At the episode level, GiGPO samples a group of complete trajectories under identical task and initial-state conditions, and computes macro relative advantages based on total returns like vanilla GRPO [18]. This captures the overall effectiveness of each trajectory and reflects the completeness of task execution. (ii) At the step level, GiGPO introduces a novel anchor state grouping mechanism for fine-grained relative advantages estimation. Specifically, it retroactively identifies the repeated environment states, termed as anchor states, across trajectories and uses them as anchors to construct step-level groups that allow for localized credit assignment.
@@
We first evaluate GiGPO on two long-horizon agent benchmarks: ALFWorld [5], which tests embodied task planning in simulated household environments, and WebShop [22], which simulates complex, goal-driven web interactions. In addition, we study multi-turn tool-integrated reasoning on search-augmented QA tasks. Our experiments with Qwen2.5-1.5B/3B/7B-Instruct [3] show that GiGPO consistently outperforms prompt-based agents, actor-critic baselines, and prior group-based RL methods. In particular, GiGPO injects fine-grained, step-level credit signals that sharpen policy learning of agents over horizons and achieves performance improvement of > 12% on ALFWorld and > 9% on WebShop over GRPO, along with remarkable performance on search-based QA tasks (42.1% on 3B and 47.2% on 7B). These gains come without compromising the core strengths of group-based RL (only < 0.002% time cost), making GiGPO a versatile and high-utility algorithm for the optimization of LLM agents.
@@ 2 Related Work
Reinforcement learning for LLM agents. RL has played a pivotal role in enabling LLM agents to operate in dynamic, open-ended environments. Early work applied classical RL algorithms such as DQN [35] to train LLM agents in text-based games [36] and later research [37, 38, 39, 40, 41] started to employ value-based methods, such PPO [42] and AWR [43], in more diverse and interactive agent scenarios including Android device control [44], embodied ALFWorld [5], and card games [45]. More recent approaches have extended RL training to complex web-based and application-centered tasks. For instance, ArCHer [46] and AgentQ [47] target the WebShop benchmark [22], but require intricate designs and computation overhead such as additional value networks or Monte Carlo Tree Search (MCTS) [48]. CoSo [10] introduces an entropy-based RL method that enhances the performance of agents. Going further, LOOP [49] introduces a hybrid method combining REINFORCE leave-one-out (RLOO) [16, 17] with PPO-style updates, achieving state-of-the-art results in AppWorld [50]. RAGEN [51] introduces a trajectory-level GRPO that concatenates all states, intermediate reasoning, and actions into a unified episode-level response. However, it faces scalability challenges in long-horizon tasks (e.g., in ALFWorld, which involves up to 50 steps).
@@ 5 Experiment
In this section, we present empirical evaluations of GiGPO across a variety of agentic tasks. Specifically, we aim to demonstrate: (1) the strong ability of GiGPO in training LLM agents; (2) the ablation study of GiGPO; (3) the dynamic trend of step-level group over the course of training; (4) the computational budget of GiGPO.
@@ 5.1 Experiment Setup
Benchmarks. We first train the LLM agents on two challenging benchmarks: ALFWorld [5] and WebShop [22]. ALFWorld is an embodied environment designed to assess the ability of LLM agents to perform multi-step decision-making. In each episode, the agent receives a text goal and must accomplish it through multi-turn interaction with the environment. It includes 3,827 task instances across six categories of common household activities: Pick & Place (Pick), Examine in Light (Look), Clean & Place (Clean), Heat & Place (Heat), Cool & Place (Cool), and Pick Two & Place (Pick2). WebShop is a complex, web-based interactive environment designed to test the LLM agents in realistic online shopping scenarios. To complete the task, the agent must interact with a simulated HTML-based shopping website to search for, navigate to, and ultimately purchase a suitable item. It contains over 1.1 million products and 12k user instructions, providing a rich and diverse action space. In addition, we also evaluate the multi-turn tool calling performance of GiGPO on search-augmented QA tasks, including single-hop QA datasets (NQ [62], TriviaQA [63], and PopQA [64]) and multi-hop QA datasets (HotpotQA [65], 2Wiki [66], MuSiQue [67], and Bamboogle [68]).
Baselines. For ALFWorld and WebShop, we compare our approach with a range of competitive baselines: (1) Closed-source LLMs: GPT-4o [1] and Gemini-2.5-Pro [2], which represent state-of-the-art capabilities in general-purpose reasoning and language understanding. (2) Prompting agents: ReAct [29] and Reflexion [30], which rely on in-context prompting to guide multi-step behavior without parameter updates. (3) RL training methods: PPO [42], a widely used actor-critic algorithm that requires an additional value model, and group-based critic-free methods RLOO [16, 17] and GRPO [18], which perform advantage estimation over trajectory groups. For search-augmented QA, we compare GiGPO with R1-Instruct, Search-R1 [58], ZeroSearch [59], and StepSearch [69].
Training details. We use Qwen2.5-1.5B/3B/7B-Instruct [3] as our base models. The weighting coefficient is set to with no further tuning. For ALFWorld and WebShop, all RL training methods (including ours and the baselines) use exactly the same hyperparameter configurations. The rollout group size for group-based RL methods is set to 8. For search-augmented QA, we follow the same settings in Search-R1 [58]. We use E5 [70] as the retriever. The rollout group size is set to 5 and the max turn is set to 4. Moreover, we incorporate similarity-based GiGPO, where anchor state grouping is performed by grouping two states if their similarity (longest matching subsequence) exceeds the threshold of 0.9. Full training settings and hyperparameter details are provided in Appendix D.1.
5.2 Performance on ALFWorld and WebShop
@@
| Closed-Source Model | ||||||||||
@@
5.3 Performance on QA tasks
| Type | Method | Single-Hop QA | Multi-Hop QA | ||||||
| NQ† | TriviaQA⋆ | PopQA⋆ | HotpotQA† | 2Wiki⋆ | MuSiQue⋆ | Bamboogle⋆ | Avg. | ||
| Qwen2.5-3B-Instruct | |||||||||
| RL Training | R1-Instruct | 27.0 | 53.7 | 19.9 | 23.7 | 29.2 | 7.2 | 29.3 | 27.1 |
| RL Training | Search-R1 | 34.1 | 54.5 | 37.8 | 32.4 | 31.9 | 10.3 | 26.4 | 32.5 |
| RL Training | ZeroSearch | 41.4 | 57.4 | 44.8 | 27.4 | 30.0 | 9.8 | 11.1 | 31.7 |
| RL Training | StepSearch | - | - | - | 34.5 | 32.0 | 17.4 | 34.4 | - |
| RL Training | GiGPO | 42.0 | 59.5 | 42.4 | 36.9 | 37.0 | 12.6 | 64.1 | 42.1 |
| Qwen2.5-7B-Instruct | |||||||||
| RL Training | R1-Instruct | 21.0 | 44.9 | 17.1 | 20.8 | 27.5 | 6.0 | 19.2 | 22.4 |
| RL Training | Search-R1 | 39.3 | 61.0 | 39.7 | 37.0 | 40.1 | 14.6 | 36.8 | 38.5 |
| RL Training | ZeroSearch | 43.6 | 61.8 | 51.5 | 34.6 | 35.2 | 18.4 | 27.8 | 39.1 |
| RL Training | StepSearch | - | - | - | 38.6 | 36.6 | 22.6 | 40.0 | - |
| RL Training | GiGPO | 46.4 | 64.7 | 46.1 | 41.6 | 43.6 | 18.9 | 68.9 | 47.2 |
As shown in Table 2, GiGPO achieves strong and consistent gains on multi-turn search-augmented QA tasks, reaching 42.1% at 3B and 47.2% at 7B, and significantly outperforming prior strong baselines such as Search-R1 and StepSearch. Although search-augmented QA is relatively short-horizon, the step-level signals of GiGPO still yield meaningful improvements. Furthermore, we observe that GiGPO is markedly more tool-efficient. Under a limit of at most 3 tool calls per query, the 7B model requires only 0.9 calls on average for single-hop tasks and 1.6 calls on average for multi-hop tasks: matching the superior performance of OTC [61], which achieves 1.0 and 1.7 calls, respectively, on the same tasks. This efficiency likely stems from GiGPO’s ability to effectively identify and suppress redundant queries in multi-turn decision-making. For instance, in repetitive patterns such as query1 info1 query1 info1 query1 info1, GiGPO groups them into the same step-level group and eliminates unnecessary repeated calls during the training, similar to the example we introduced in Figure 3.
@@ 6 Conclusions and Limitations
In this work, we proposed GiGPO, a novel group-based RL algorithm to tackle the credit assignment challenge in long-horizon LLM agent training. GiGPO introduces a hierarchical advantage estimation that enables fine-grained per-step credit assignment while retaining the efficiency and stability of group-based RL. By retroactively grouping steps that share the same state across trajectories, it achieves this without incurring additional LLM rollout or GPU memory overhead. Empirical evaluations across complex agentic environments (ALFWorld and WebShop) and search-augmented QA tasks demonstrate that GiGPO significantly outperforms both prompt-based agents and prior RL methods. A potential limitation of GiGPO is its reliance on state matching for anchor group construction. In highly complex environments, identical states may be hard to detect due to noise or subtle differences. Despite this, GiGPO still retains a strong performance lower bound: in the extreme case where no states are repeated across trajectories (i.e., ), it naturally degrades to GRPO, preserving GRPO’s effectiveness and stability in credit assignment. Although this issue is partly mitigated by incorporating similarity-based grouping, exploring more robust state-matching strategies, such as embedding-based representations or domain-specific structural equivalence, remains an important direction.
@@ Appendix A Source Code
Our code is built upon the veRL framework [73] and extended with several features to support scalable training of long-horizon LLM agents. Key capabilities of our code include: (1) a step-wise multi-turn interaction paradigm that avoids concatenating the entire interaction history into a single input, as done in RAGEN [51], thereby ensuring scalability for very long-horizon optimization, (2) support for parallelized, gym-style group-based environments to enable high-throughput and state-aligned rollouts, (3) a rich suite of environments, including Search (tool use), ALFWorld, WebShop, Sokoban, and Gym Cards; and (4) compatibility with both text-based and vision-language agents.
@@ D.1 Details of Training
Hyperparameters for Search-Augmented QA. The maximum prompt length is 4096 tokens, and the maximum response length is 512 tokens. The max turn is set to 4. The learning rate is 1e-6 for the actor. We adopt a rule-based reward, assigning a reward of 1 for success and 0 for failure. Invalid actions are penalized with a reward of -0.01. We set the train data size to 256 and use a group size of 5. Rollout and validation temperatures are set to 1.0 and 0.0, respectively. The mini-batch size is 512, and the KL-divergence loss coefficient is set to 0.001. The weighting coefficient is set to 1 without additional tuning, and the discount factor is set to 0.95.
Computing Details. For ALFWorld and WebShop, Qwen2.5-1.5B experiments are run on 2×H100 GPUs and Qwen2.5-7B on 4×H100 GPUs, each for 150 iterations. For search-augmented QA, Qwen2.5-3B uses 4×H100 GPUs and Qwen2.5-7B uses 8×H100 GPUs, each for 200 iterations.
@@ D.2 Prompts
The prompts we use for LLM agents are presented in Figure 7, Figure 8, and Figure 9. These prompt templates are constructed using Python-style string formatting, where placeholders enclosed in curly braces ({}) represent semantic slots. These placeholders, such as {task_description}, {step_count}, and {current_observation}, are dynamically populated at runtime via Python’s .format() function. To enrich the agent’s context, we use historical information and set the history length to 2 for ALFWorld and WebShop and the full history for search-augmented QA experiments.
The <think> </think> block instructs the agent to explicitly perform step-by-step reasoning, thereby promoting chain-of-thought [74] style deliberation. The <action> </action> block is used to clearly indicate the final action decision. The search agent outputs reasoning traces within <think> </think>, issues search queries within <search> </search>, provides anwsers within <anwser> </anwser>. Retrieved evidence from the retriever is presented in <information> </information> tags.